can you please help me to solve this equation?
400r=x^2-40x+400
i want the answer in r

Answer by Alan3354(69443)

400r=x^2-40x+400
r = (x^2-40x+400)/400
-----------------
Like that?
-----------
x^2-40x+400(1-r) = 0
x = 20 ± 20*sqrt(r) or like that?